Wed 31 Oct 2007 - Most systems ran without problems
Between 15:00 to 16:30, Outlook Web Access, POP and IMAP services were unavailable.
Tue 30 Oct 2007 - Most systems ran without problems
The scheduled PC lab upgrade work in Engineering (including George Fox and Whewell buildings) was completed without problem.
Mon 29 Oct 2007 - All systems ran without problems
Fri 26 Oct 2007 - All systems ran without problems
Departmental filestore was returned to normal service by 7:00 PM Thursday 25th October
Personal filestore (h:drive) was returned to normal service by 12:10 AM Friday 26th October
Any office machines that are behaving in an erratic manner after 7am Friday 26th October should be logged of and logged back on again which should resolve any connection issues with the central file servers.
Thu 25 Oct 2007 - Central Filestore Maintenance
The central file servers that provide personal stores (h: drive) and departmental storage were taken out of service for essential planned maintenance from 6:00 PM Thursday 25th October 2007 and were set to be returned to service by 7:00 AM Friday 26th October.
Due to unforeseen hardware failure on the central file servers we believe we have suffered some minor file corruption, this has been rectified. In order to prevent further corruption and potential data loss ISS are taking a proactive action to check and fix any errors on the central file stores. We are performing this maintenance overnight due to the length of time required to perform these operations which would be prohibitive during the working day.
Managed XP PC build users should be unaffected by this process however PC labs and 'non-managed machines' may operate at a reduced level of functionality during this period.
Wed 24 Oct 2007 - Exchange Maintenance - Most systems ran without problems
The Exchange Email Systems will be intermittently unavailable between 1PM and 4PM on Wednesday 24th October, for the application of security updates. We apologise for any inconvenience caused.
Tue 23 Oct 2007 - Most systems ran without problems
Some problems with Webmail early in the day. These were caused by unxepected demands on the mail server.
Mon 22 Oct 2007 - Most systems ran without problems
Fri 19 Oct 2007 - Most systems ran without problems
The second batch of Exchange users are being moved back on to the repaired server.
Thu 18 Oct 2007 - Most systems ran without problems
Phones on the Voice over IP system are currently having problems in the LEC area. ISS are investigating the issue and will restore service as soon as possible.
The first batch of Exchange users are being moved back on to the repaired server.
The server that had hardware problems has been repaired and tested. When the server failed moved users to two other servers. These moved users are going to be moved back on to the repaired server on Thursday 18th Oct and Friday 19th Oct. Affected users Users can expect to see the “the administrator has made a change which requires you to restart outlook.” Users should then restart Outlook.
Wed 17 Oct 2007 - Most systems ran without problems
Printers in the Library are still not reporting paper-out and are being manually checked. The problem is understood and will be resolved in due course.
Tue 16 Oct 2007 - Most systems ran without problems
Some network printers have stopped reporting their 'paper-out' status. Network printers are currently being manually inspected for 'paper-out' status and the cause of the problem is being investigated.
Mon 15 Oct 2007 - All systems ran without problems
Fri 12 Oct 2007 - All systems ran without problems
Thu 11 Oct 2007 - All systems ran without problems
Wed 10 Oct 2007 - Most systems ran without problems
LUVLE, My Modules, MyPlace, Online Assessments and Module Feedback will be unavailable between 2pm and 4pm to allow for essential server maintenance work to be carried out.
Tue 09 Oct 2007 - All systems ran without problems
Mon 08 Oct 2007 - 'Homes 3' On Central Filestore Unavailable Until 10:10
Home server 3 (folders 48-60) was unavailable. Some users were unable to access H Drive. The problem was investigated and the full filestore service restored at 10:10.
Fri 05 Oct 2007 - Most systems ran without problems
Thu 04 Oct 2007 - 'Homes2' On Central Filestore Unavailable
The 'Homes2' area of central filestore has suffered a hardware failure and is currently unavailable. Users with folder numbers between 20 & 39 are on Homes2 and affected. ISS technicians are working to restore the service. The service was resolved at 12:45 without data loss.
Wed 03 Oct 2007 - All systems ran without problems
Equipment serving the Voice Over IP telephone network on campus will be taken down for essential maintenance tonight at 6:00pm. If your extension number begins 10xxx then you will be affected by this work. The work will not affect calls that are in progress, but you may notice your phone reconfigure itself for a period of about 30 seconds during the maintenance window. Once your phone has reconfigured you will be able to make calls again. Work will begin at 6:00pm and should be complete by 7:00pm. The completion time will be posted on Message of the Day.
Please note that whilst phones are reconfiguring 999 calls will be unavailable.
High Performance Cluster
The HPC head node corona is now available again to users, and the cluster is fully back in service.
As the fix required a reboot of all nodes, a small number of still-running jobs were terminated - after logging on to the head node,
type: cat /tmp/lost_jobs for a list.
Tue 02 Oct 2007 - Most systems ran without problems
During the Network 'at risk' period on 02/10/07 work was successfully carried out on core devices of WiNet.
During the 'at risk' period on 02/10/07 work was successfully carried out on a core network device in Infolab.
Mon 01 Oct 2007 - Most systems ran without problems
The HPC will remain unavailable today due to the ongoing systems problem. The cluster vendor, Streamline Computing, are currently working on a fix, and hope to provide a solution by close of business today, with a view to our applying the fix to the HPC tomorrow.
Over the weekend the 'ResNet Validation' system successfully validated 'all comers' with two short breaks in service.